an aqueous solution is 4.30m lithium chloride. the density of the solution is 1.127 g/mL calculate a) molarity b) mole fraction c) percent by mass
Answer #1

molality = 4.30m

molality is the number of moles of solute present in a 1 Kg of a solvent

moles of lithium chloride = 4.30 mol

mass of lithium chloride present in solution = 4.3 * 42.394 = 182.29 g

mass of water = 1000 g

mass of a solution = 1000 + 182.29 = 1182.29

density = mass / volume

volume = mass / density

volume = 1182.29 / 1.127

volume of a solution = 1049.059 mL

a)

molarity = moles / volume of a solution in liter

molarity = (4.3 / 1049.059 ) *1000

molarity = 4.098 M

b)

moles of water = 1000 / 18 = 55.55 mol

moles fraction of lithium chloride = 4.3 / 4.3 + 55.55

= 4.3 / 59.85

moles fraction of lithium chloride =  0.0718  

moles fraction of water = 1 - 0.0718   = 0.9282

c)

percentage by mass = ( mass of a compound / total mass of a solution ) * 100

= (182.29 / 1182.29 ) *100

percentage by mass = 15.42 %